OpenStudy (anonymous):

Which of the following equations has this graph? Answer y = -1/2 x + 3 y = -2x +3 y = 3x - 2 y = 1/2x - 3

OpenStudy (anonymous):

very confused

sam (.sam.):

The first thing you can tell it is either A or B because since the slope is  And the slope for both y = -1/2 x + 3 y = -2x +3 Is negative

OpenStudy (anonymous):

First of all, the graph is decreasing, so the value that multiplies x have to be negative. f(0) = 3. You now can have the two first functions. But the root of the equation is 6. So, it's the first equation: -1/2x + 3.

sam (.sam.):

y=mx+c m is the slope

OpenStudy (anonymous):

okay so once you have m what do you do Im still lost

sam (.sam.):

Just use a coordinate from the graph, example (0,3) and (6,0) Find the slope, then use the point slope formula.

sam (.sam.):

Slope formula is \[Slope=\frac{y_2-y_1}{x_2-x_1}\] Substitute and find slope You should get -1/2, then you'll know it's y = -1/2 x + 3

OpenStudy (anonymous):

thanks